Chelsi was crossing a bridge constructed on a river. While standing on the bridge, she saw 2 boats Liberty and Columbus cross from the opposite side of the river. She could see liberty at an angle of depression of 60° and Columbus at 45°.Chelsi’ s height is 168cm and the height of the bridge from the ground is 170 m from the ground. Two people riding the boats saw the opposite edge of the bridge at an angle of 45. Find the distance between the 2 boats and also the distance to cover by the two boats to reach exactly below the bridge where Chelsi is standing.

Grade:9

1 Answers

Ankit
6 years ago
The diagram should be clear to u....Let AD be the total height from the ground including. Chelsi s height ........... AD =171.68m...............A triangle ACD is formed on the side of liberty where angle of depression is 60°. ....... Therefore angle ACD is also 60°.. . By applying tan60 we get CD the distance between boat and exact point at the ground below chelsi...........CD= 171/√3M................ In the same way on colombus side applying tan 45 we get the distance DE in Triangle. ADE which is 171 m......................... By adding them we get distance between 2 boats that is 171.68/√3+ 171.68m. = 171.68(1/√3+1)m...........